OpenStudy (anonymous):

If f(x) = 5x – 3 and g(x) = 3x – 3, find f(x) – g(x). A. 2x B. 8x – 6 C. 2x – 6 D. 8x

OpenStudy (anonymous):

You just have to subtract the equations. f(x) – g(x) = (5x – 3) -(3x – 3) Now you just have to simplify

OpenStudy (anonymous):

(f-g)(x) is a shorthand way of saying that you need to get the difference of the two functions: f(x) and g(x). so essentially, (f-g)(x) = f(x) - g(x) = 5x+2 - (-3x+4) = 8x-2 (f/g)(-2) = f(-2)/g(-2) = -4/5 or -0.8 Hope this helps.